# A Benchmark for Multi-speaker Anonymization

### v1

@@ VI-A1 Multi-speaker anonymization configurations
For SD, we apply the efficient and robust spectrum clustering-based approach, which is implemented with the WeSpeaker Toolkit  [^6] https://github.com/wenet-e2e/wespeaker/tree/master [49]. It first performs Silero-VAD  [^7] https://github.com/snakers4/silero-vad to remove the silent segments, then splits the long audio into the 1.5-second segments. Note that these segments have 0.75 seconds of overlapping. Each segment is fed into the speaker recognition model, which is a context-aware masking-based structure [50] pre-trained on the VoxCeleb2 dataset [51]. After that, spectral clustering [39, 40] is performed to aggregate the segments into several speakers by analyzing the similarity metrics. The diarization results are saved into the RTTM file, which summarizes the timestamp of each speaker’s active speech segments.
@@
$A_{Select}$, the commonly-used speaker anonymizer, uses a selection-based strategy that identifies the 200 farthest same-gender speaker vectors from an external speaker vector pool (specifically LibriTTS-train-other-500 [55]), subsequently averaging 10 randomly-selected ones as the pseudo-speaker vector99910 vectors instead of the default setting of 100 used in VPC, as it has been demonstrated that averaging 100 (large number) vectors significantly reduces the distinctiveness of anonymized speakers [17]..
@@ VI-C2 Results on simulated datasets
The first row of Table VII shows the WERs on the original LibriSpeech test-clean subset used to simulate the conversation and its anonymized audios using different speaker vector anonymization methods. The second row presents the results for the noisy dataset, which adds background noises to the original LibriSpeech test-clean subset. All anonymized audios obtain higher WERs than those of original audios. For the clean condition, the absolute difference between the original (1.89%) and different anonymized audios (around 2.50%) is about 0.61%. For the noisy condition, this difference increases to about $7.12\%=13.70\%-6.58\%$, as $ASR_{\text{eval}}$ is trained on clean speech, leading to more mismatches when decoding noisy speech. The differences among different anonymization methods are minor and the speech content is preserved at an acceptable level after anonymization.
@@
In summary, the proposed MSA systems using $A_{DS}$ and $A_{AS}$ achieve the overall best performance in terms of various privacy and utility metrics.
@@ VI-C3 Results on VoxConverse datasets
VI-D Discussion on overlapping speech
Until now, we have verified the effectiveness of proposed MSA systems on non-overlapping conversations. It is interesting to explore scenarios where overlapping segments exist, with a specific focus on the privacy risks to speakers in adjacent overlapping segments—both preceding and succeeding speakers. We consider the same attacker as in non-overlapping scenarios who uses $ASV_{eval}$ to infer the adjacent speaker identities from the anonymized overlapping segments [^16] Note that a stronger attacker may infer the original speaker’s identity after separating overlapping segments, we leave it as a future work.. Specifically, we select 97 conversations from VoxConverse with overlaps and attempt to answer the following three questions. Does the overlapping segment reveal the privacy of nearby single speakers? We crop overlapping segments from the 97 selected conversations and extract the connected preceding and following single-speaker segments. Overlapping segments and either preceding or following single-speaker segments are taken as negative pairs. The single-speaker segments, split in half, are taken as positive pairs, as shown in Fig. 7a. Fig. 7b plots the cosine similarity between overlapping segments and their nearby single-speaker segments, including preceding and following single-speaker segments (blue), as well as the cosine similarity for segments from the same speaker (orange). We treated the former as negative pairs and the latter as positive pairs to compute the EER. Ideally, an EER of 0% indicates the overlapping segments do not leak adjacent speaker identity information. However, an EER of 13.40% was obtained, suggesting a low level of privacy leakage.
@@ VI-D Discussion on overlapping speech
What can be done to avoid privacy leakage from overlapping segments? One intuitive approach is to detect overlapping segments, apply speech separation to them, anonymize, and then reconstruct. However, this method is resource-intensive. Considering human speaking habits, overlapping segments typically do not convey too much important information. If something important is said, the speaker usually repeats it. Thus, one quick solution is to detect the overlap region and remove it entirely. Another method is to detect the overlap region and shuffle the segments along the time axis. Fig. 7d plots the cosine similarity, which is similar to Figure 7b, but with the overlapping segments shuffled along the timestamp. This shuffle reduces the EER from 13.40% to 1.98% as shown in Fig. 7d.
VII Conclusion and Future Work
@@
In this study, we consider that the attacker does not have knowledge or access to the MSA system and use an ASV evaluation model to identify the original speaker identity from anonymized speech. In addition, in the overlapping discussion in Section VI-D, we assume the attacker infers nearby speaker identities from overlapping segments, rather than separating overlapping segments first and then inferring original speakers. However, in real-life scenarios, a stronger attacker might possess prior knowledge of the MSA system, as well as the ability to separate overlapping segments and identify original speakers. Thus, in future work, we will investigate an end-to-end MSA approach that can handle both non-overlapping and overlapping conversations simultaneously, with an uncertain number of speakers, and evaluate its privacy protection abilities under various attack scenarios.

### v2

@@ VI-A1 Multi-speaker anonymization configurations
For SD, we apply an efficient and robust spectrum clustering-based approach, which is implemented with the WeSpeaker Toolkit  [^9] https://github.com/wenet-e2e/wespeaker/tree/master [54]. It first uses Silero-VAD  [^10] https://github.com/snakers4/silero-vad to remove silent segments, then, we follow the common setting from [45] and split the long audio into 1.5-second segments with a 0.75-second overlap. Each segment is fed into the speaker recognition model, which uses a context-aware masking-based structure [55] pre-trained on the VoxCeleb2 dataset [56]. After that, spectral clustering [43, 44] is performed to aggregate the segments into several speakers by analyzing similarity metrics. The diarization results are saved into the RTTM file, which summarizes the timestamps of each speaker’s active speech segments.
@@
$A_{Select}$, the commonly-used speaker anonymizer, uses a selection-based strategy that identifies the 200 farthest same-gender speaker vectors from an external speaker vector pool (specifically LibriTTS-train-other-500 [60]), subsequently averaging 10 randomly-selected ones as the pseudo-speaker vector12121210 vectors instead of the default setting of 100 used in VPC, as it has been demonstrated that averaging 100 (large number) vectors significantly reduces the distinctiveness of anonymized speakers [20]..
@@ VI-C2 Results on simulated datasets
The first row of Table VII shows the WERs on the original LibriSpeech test-clean subset used to simulate the conversation and its anonymized audios using different speaker vector anonymization methods. The second row presents the results for the noisy dataset, which adds background noises to the original LibriSpeech test-clean subset. All anonymized audios obtain higher WERs than those of original audios. For the clean condition, the absolute difference between the original (1.89%) and different anonymized audios (around 2.50%) is about 0.61%. For the noisy condition, this difference increases to about $8.12\%=13.70\%-5.58\%$, as $ASR_{\text{eval}}$ is trained on clean speech, leading to more mismatches when decoding noisy speech. The differences among different anonymization methods are minor and the speech content is preserved at an acceptable level after anonymization.
@@
In summary, the proposed MSA systems using $A_{{DS}}$ and $A_{{AS}}$ consistently achieve the best performance in terms of DER. For WER and PMOS, $A_{Select}$, $A_{{AS}}$, and $A_{{DS}}$ are comparable, with most cases showing slightly worse performance for $A_{{AS}}$ and $A_{{DS}}$ compared with $A_{Select}$. For FAR, when using real RTTM under clean conditions, $A_{Select}$ achieves the lowest FAR, and the differences between $A_{Select}$ and $A_{{DS}}$, $A_{{AS}}$ are the largest. However, these differences become progressively smaller under noisy conditions when using predicted RTTM, or with a larger number of speakers. This indicates that $A_{{AS}}$ and $A_{{DS}}$ are more robust.
@@ VI-C3 Results on VoxConverse datasets
VII Discussions
VII-A Discussion on overlapping speech
Until now, we have verified the effectiveness of proposed MSA systems on non-overlapping conversations. It is interesting to explore scenarios where overlapping segments exist, i.e., a cocktail party scenario, with a specific focus on the privacy risks to speakers in adjacent overlapping segments—both preceding and succeeding speakers.
We consider the same attacker as in non-overlapping scenarios who uses $ASV_{eval}$ to infer the adjacent speaker identities from the anonymized overlapping segments [^19] Note that a stronger attacker may infer the original speaker’s identity after separating overlapping segments, we leave it as a future work.. Specifically, we select 97 conversations from VoxConverse with overlaps and attempt to answer the following three questions.
Does the overlapping segment reveal the privacy of nearby single speakers? We crop overlapping segments from the 97 selected conversations and extract the connected preceding and following single-speaker segments. Overlapping segments and either preceding or following single-speaker segments are taken as negative pairs. The single-speaker segments, split in half, are taken as positive pairs, as shown in Fig. 8a. Fig. 8b plots the cosine similarity between overlapping segments and their nearby single-speaker segments, including preceding and following single-speaker segments (blue), as well as the cosine similarity for segments from the same speaker (orange). We treated the former as negative pairs and the latter as positive pairs to compute the EER. Ideally, an EER of 0% indicates the overlapping segments do not leak adjacent speaker identity information. However, an EER of 13.40% was obtained, suggesting a low level of privacy leakage.
@@ VII-A Discussion on overlapping speech
What can be done to avoid privacy leakage from overlapping segments? A naive approach is to sacrifice the utility when processing the overlapped segments. This can be done by detecting and shuffling the segments along the time axis. Fig. 8d plots the cosine similarity, which is similar to Figure 8b, but with the overlapping segments shuffled along the timestamp. This shuffle reduces the EER from 13.40% to 1.98% as shown in Fig. 8d. Despite the decreased EER, the naive method severely degrades the linguistic information in the overlapped segments. A potentially better pipeline may apply speech separation to the overlapped region, anonymize the separated segments, and reconstruct the utterance. This work raises awareness of the overlapping segments; however, the implementation and evaluation of more advanced pipelines are left for future work.
VII-B Discussion on evaluation metrics
In this paper, due to the high cost of recruiting human subjects for listening tests, we rely on objective metrics, including FAR, WER, PMOS, and DER. However, a key question arises: can objective evaluation metrics fully capture the subjective aspects of speech? To assess the naturalness of anonymized speech, we use a MOS prediction network to compute PMOS. Previous work [31] demonstrated a similar ranking of speaker anonymization systems between human-rated MOS scores and those predicted by the MOS network, suggesting that the network’s predictions align well with human perception. In addition, we use WER to evaluate intelligibility. Although the VoicePrivacy 2022 summary paper [69] highlights that the correlation between WER and intelligibility is not particularly strong, it emphasizes that both objective and subjective metrics provide valuable insights.
Figure 9: Privacy-to-utility trade-off for different MSA systems on clean simulated conversations, where FAR and DER are averaged across different numbers of speakers and calculated using the predicted RTTM. Evaluations are conducted for $\lambda\in[0.1,0.3,0.5,0.7,0.9]$.
Another limitation of the current metric computation is that the use of multiple evaluation metrics complicates the assessment process and makes comparison difficult, particularly because of the trade-off between privacy and utility, e.g. $A_{Select}$ achieves a lower FAR at the expense of conversational dynamics degradation, yielding a higher DER. To address this, we propose the privacy-to-utility trade-off, a compressed metric that combines the primary privacy metric and the primary utility scores, with a weight to control the importance of privacy and utility. Inspired by this, we adopt a similar compressed metric, privacy-to-utility trade-off ($\mathrm{PU_{tr}}$), which combines the primary privacy metric (FAR) and the primary utility scores (WER, DER, PMOS), designed to evaluate the anonymization system at different operating points. Given $\mathrm{WER_{i}},\mathrm{DER_{i}},\mathrm{PMOS_{i}},\mathrm{FAR_{i}},i\in\{0,1\}$ denotes the metrics calculated on original ($i=0$) or anonymized ($i=1$) utterances.
$\mathrm{PU_{tr}}=$ $\ {\lambda}\left(\frac{\log\left(1+\frac{\mathrm{WER_{1}}}{\mathrm{WER_{0}}}\right)}{\log\left(1+\frac{1}{\mathrm{WER_{0}}}\right)}+\frac{\log\left(1+\frac{\mathrm{DER_{1}}}{\mathrm{DER_{0}}}\right)}{\log\left(1+\frac{1}{\mathrm{DER_{0}}}\right)}-\frac{\log\left(1+\frac{\mathrm{PMOS_{1}}}{\mathrm{PMOS_{0}}}\right)}{\log\left(1+\frac{1}{\mathrm{PMOS_{0}}}\right)}\right)$
$+(1-\lambda)\cdot\frac{\log\left(1+\frac{\mathrm{FAR_{1}}}{\mathrm{FAR_{0}}}\right)}{\log\left(1+\frac{1}{\mathrm{FAR_{0}}}\right)}$
where $\lambda\in[0,1]$ controls the trade-off between utility and privacy. A lower $\mathrm{PU_{tr}}$ indicates a better trade-off at a specific operational point $\lambda$. Figure 9 shows the $PU_{tr}$ results for different MSA systems on clean simulated conversations, where FAR and DER are averaged across different numbers of speakers and calculated using the predicted RTTM. Lower $\lambda$ prioritize privacy, while higher $\lambda$ emphasize utility. It is clear that as $\lambda$ increases, MSA using $A_{AS}$ and $A_{DS}$ achieves a better trade-off than other MSA systems, particularly when $\lambda>0.5$. Within $A_{AS}$ and $A_{DS}$, $A_{AS}$ shows a better trade-off. This further confirms the effectiveness of $A_{DS}$ and $A_{AS}$, particularly in utility-DER.
VII-C Discussion on MSA computational efficiency
We have evaluated the effectiveness of the proposed MSA in terms of privacy and utility performance. However, it is equally important to analyze the latency, computational demands, and practicality of the MSA for real-world, latency-sensitive applications. The system achieves an average Real-Time Factor (RTF) of 0.29, which represents the ratio of the total processing time to 5 hours of audio [^20] Simulations were conducted in both clean and noisy audio environments, involving 2–5 speakers, each contributing 5 hours of audio. The total processing time for eight different simulation datasets ranged from 4,444 to 7,057 seconds., as tested on an NVIDIA A5000 GPU, indicating that the system operates approximately three times faster than real-time. In terms of computational demands, the multispeaker anonymization system comprises approximately 130 million (130M) parameters, distributed among its key components: speaker diarization (15M), the HiFi-GAN Generator (14M), the HuBERT-soft Encoder (95M), and the ECAPA-TDNN (6M). Generating 540 seconds of audio with these models requires approximately 40 GB of RAM, making it costly and challenging to scale or deploy on edge devices. One possible solution, as proposed recently, is to replace traditional non-causal, computationally intensive networks (e.g., HuBERT and HiFi-GAN) with lightweight convolutional neural network architectures, thereby achieving low latency. Further optimization strategies are left for future research. One possible solution, as proposed recently by [70], is to replace traditional non-causal, computationally intensive networks (e.g., HuBERT and HiFi-GAN) with lightweight convolutional neural network architectures, thereby achieving low latency. Further optimization strategies are left for future research.
VII-D Discussion on SSA and MSA computational efficiency and privacy-utility trade-offs
It is evident that SSA is less resource-intensive, as it does not require speaker diarization or conversation-level speaker anonymizers. By directly applying SSA to multi-speaker anonymization, privacy is simultaneously enhanced, as all speakers are converted into a single speaker, making it difficult to identify the source speakers. However, this approach completely erases multi-speaker relationships, significantly reducing utility. Overall, SSA and MSA present different trade-offs: while SSA offers a more computationally efficient solution with greater privacy, it sacrifices utility by removing speaker-specific interactions. Conversely, MSA maintains multi-speaker relationships, offering higher utility but requiring more computational resources. The choice between SSA and MSA depends on the specific application and the relative importance of privacy and utility.
VIII Conclusion
@@
Future work could focus on: (i) Developing an end-to-end real-time MSA approach for overlapping and non-overlapping conversations: This paper presents a multi-speaker anonymization benchmark using a cascaded system, with our current focus primarily on non-overlapping conversations. However, overlapping speech segments also have the potential to reveal source speaker information and pose unique challenges for anonymization. Future work could investigate strategies to handle overlapping conversations effectively, even with an uncertain number of speakers. Additionally, improving the system’s efficiency in terms of latency and computational demands is critical to enabling real-time applications, making this another important direction for future research. (ii) Enhancing evaluation metrics for a more comprehensive analysis: Although the proposed metric $\mathrm{PU_{tr}}$ is a novel attempt to integrate multiple objective evaluation metrics into a single measure, it relies on a hyperparameter (the weight $\lambda$) to balance privacy and utility. Future work could focus on refining the evaluation framework to streamline the assessment process, potentially by incorporating both subjective (e.g., human perceptual studies) and objective metrics. This would provide a more comprehensive analysis of the naturalness and practical utility of anonymized speech, addressing the limitations of the current approach.